OpenStudy (anonymous):

Determine the zeros of f(x) = x3 + 7x2 + 10x − 6

OpenStudy (anonymous):

Im lost :(

ganeshie8 (ganeshie8):

familiar with rational root theorem ?

OpenStudy (anonymous):

I think so

ganeshie8 (ganeshie8):

good, what does it say ?

OpenStudy (anonymous):

A theorem that provides a complete list of possible rational roots of the polynomial equation

OpenStudy (anonymous):

Would it look like this give me a min

OpenStudy (anonymous):

f(x) = 1x*3 + 7x*2 + 10x - 6 Factors of P = plus minus 1 , plus minus 2 and plus minus 3 and 6 Factors of Q = plus minus 1 P/Q = plusminus 1/1

ganeshie8 (ganeshie8):

thats right ! so the possible rational zeroes are : \[\large \pm 1 , ~\pm2, ~\pm 3, ~\pm 6\]

OpenStudy (anonymous):

Solution would be 0 = x^3 + 7x^2 + 10x - 6 is x= 0.45 and x =-4.45

ganeshie8 (ganeshie8):

plugin x = 1 in the function, do you get 0 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So there's a total of 3 zeros Positive are 1 and 1 Negative is 2 and 0 And complez is 0 and 2

OpenStudy (anonymous):

Am i correct ?

ganeshie8 (ganeshie8):

f(x) = x^3 + 7x^2 + 10x - 6 f(1) = 1^3 + 7(1)^2 + 10(1) - 6 = 1 + 7 + 10 - 6 = 12 which is NOT 0, so 1 is NOT a zero

ganeshie8 (ganeshie8):

lets try if -1 is a zero : f(x) = x^3 + 7x^2 + 10x - 6 f(-1) = (-1)^3 + 7(-1)^2 + 10(-1) - 6 = -1 + 7 - 10 - 6 = -10 which is NOT 0, so -1 is NOT a zero

ganeshie8 (ganeshie8):

similalry can you test +-2 and +-3 also ?

OpenStudy (anonymous):

Ok

OpenStudy (carolina4567):

So 2 / 1 7 10 -6 You would get 1 , 9 , 28 ,50

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Then 3 is 3 / 1 , 7 , 10 , -6 Answers would be 1 , 10 , 40 , 114 am i correct

ganeshie8 (ganeshie8):

Oh nice, you're using synthetic division is it ?

ganeshie8 (ganeshie8):

try -3 also

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

-3 / 1, 7, 10, -6

OpenStudy (anonymous):

1 , 10 , 40 ,114

ganeshie8 (ganeshie8):

thats for +3 right ?

ganeshie8 (ganeshie8):

what do you get for -3 ?

ganeshie8 (ganeshie8):

-3 / 1, 7, 10, -6 1, 4, -2, 0

ganeshie8 (ganeshie8):

so looks like -3 is a zero ?

ganeshie8 (ganeshie8):

and the depressed quadratic would be : \[\large x^2 + 4x - 2\] which you can solve using quadratic formula to find the remaining two zeroes

OpenStudy (anonymous):

OH ok i got the answer THANK U SO MUCH UR A LIFESAVER :) !!!!!!